UT Sep 29 2022 - Split night with PETS-PEPSI




Night Info

  • Sunset 18:08
  • Nautical Twilight Ends 19:01
  • Astro Twilight Ends 19:30
  • Handover to PEPSI-PETS 22:30
  • PEPSI 23:00-05:29
  • Astro Twilight Begins 04:51
  • Nautical Twilight Begins 05:19
  • Sunrise 06:12
  • Moonset 20:08
  • Moon illumination ~10%

Staff

  • OSA: J.Williams
  • ISp: C. Bolyard
  • Mtn Manager: R. Hannsen
  • ISA/SO: A.Becker, B.Rothberg, R.T. Gatto
  • AO: ?

Night Plan

ALTA prediction for Tonight

20220928_see_evol_time_Height_20_20000.png

Starting Config Bino-LUCIs

Twilight OSCO

Field Aberration Confirmatory Collection

  • Conifrmatory Field Aberration - Critical (1hour) - update installed already
  • need both AGW1 and AGW2 on the telescope and checked to execute - DOES NOT NEED LUCIs - Just AGW 1 and AGW 2
  • 1.2" or better seeing (1" or better ideal)
  • Have OSA go to field closeby, point and collimate, then preset to target
  • Target Fields to use (one):
    • BS9169_short_field.list 17 13 44 +54 33 17 - 7 stars in list - good for short confirmation
    • BS9177 18:39:33.7 +49:05:38.0
    • StoneP1 22:37:45.0 -01:13:40.45
    • BS9188 23:30:33.4 +38:18:57.0

AO On-Sky Checks

  • File: Service_Observing/Restart_2022/LUCI/AO/AO_Testing_V96.xml
  • Thin clouds, seeing 1.2" or better
  • Bin 3
    • RA 20hr-22hr folder
    • Star 6, 7, or 8 - pick a PA where the guidestars are not in the patrol field on large offsets
  • Bin 4
    • RA 20hr-22hr folder
    • Star 9, 10, or 11 - pick a PA where the guidestars are not in the patrol field on large offsets
  • Bin 2
    • RA 20hr-22hr folder
    • Star 4 or Star 5 - pick a PA where the guidestars are not in the patrol field on large offsets
  • Bin 1 (can skip if OSCO was successful, although something fainter than R~7.5 would be useful)
    • RA 20hr-22hr folder
    • Star 1 or Star 2 (star 3 needs an updated OT file for sub-frame)

ESM On-Sky Checks

  • File (mixed mode already created) - /home/lbto/Service_Observing/Restart_2022/LUCI/ESM/M52
  • M52 - mixed mode - Bin 1
  • M52 - mixed mode - Bin 4

Poor Weather Backup

  • ESM on-sky checks - M52
    • mixed mode scripts and L1 only already in place (see above)
  • Photometric Standards (LUCI) - only if clear - any seeing
    • /home/lbto/Service_Observing/2022B_D_E/PoorWeather/LUCI_Phot_ZP_OT_V6.xml
  • G200 Test on-sky
    • any conditions
    • PKS 2247+14 - bright QSO point source
  • Imaging Check
    • large dithers - pick target from OT - NGC 40 - large dithers


Night Log

  • UT 00:14 - brought up LUCI GUIs - will let Alex do the fieldstop checks
    • L1 0001-0003
  • OCAM problem? quadrant bias issue?
  • Can't run LUCI GUIs on TO1- connection error to host and unable to connect journalizer - G.Eychaner looking into it
    • G.Eychaner fixed the problem
  • UT 01:08 - opening enclosure
Googly Doc link

-- BarryRothberg - 28 Sep 2022
Topic revision: r5 - 29 Sep 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback